Use the following questions to determine, a) The cash cycle, and, b) The cash turnover rate

Sales $ 120 000
Cash sales $ 60 000
Opening stock $ 12 000
Closing stock $ 10 000
Purchase (on credit) $ 62 000
Debtors $ 5 000
Creditors $ 10 000
Fixed assets $ 65 000
Tax rate $ 25
1
Expert's answer
2012-09-06T11:51:39-0400
a) CCC = (12000 - 10000)/(60000/365) + 10000/(60000/365) - 5000/(120000/365) = (2000 + 10000 - 2500)/(60000/365) = 9500*365/60000 = 95*365/600 = 57.8 days
b) The cash turnover rate = Sales/Cash sales = $120 000/$60 000 = 2